OpenStudy (anonymous):

Multiply: (x + 1)(x - 2).

mathslover (mathslover):

\[(x+1)(x-2)\] \[x(x-2)+1(x-2)\] \[x(x)-x(2)+1(x)-1(2)\] \[x^2-2x+x-2\] \[x^2-x-2\]

jhonyy9 (jhonyy9):

x^2 -2x+x-2 =x^2 -x-2

OpenStudy (anonymous):

=x^2-x-2